# Po-Heng Chou1, Chiapin Wang2, Kuan-Hao Chen2, and Wei-Chen Hsiao2

### v2

@@ Abstract
In this paper, we propose a deep reinforcement learning (DRL)-based beam weighting framework that couples a policy network with an augmented weighted least squares (WLS) estimator for accurate and low-complexity positioning in multi-beam LEO constellations. Unlike conventional geometry or channel state information (CSI)-dependent approaches, the policy learns directly from uplink pilot responses and geometry features, enabling robust localization without explicit CSI estimation. An augmented WLS jointly estimates position and receiver clock bias, improving numerical stability under dynamic beam geometry. Across representative scenarios, the proposed method reduces the mean positioning error by 99.3% compared with the geometry-based baseline, achieving 0.395 m RMSE with near real-time inference.
@@ I Introduction
Despite these advances, most existing solutions still rely heavily on explicit channel state information (CSI) or involve computationally intensive optimization, limiting real-time applicability for resource-constrained LEO platforms. Traditional statistical estimation methods, such as weighted least squares (WLS) [10], widely used in positioning problems, provide reliable accuracy but lack adaptability in dynamic multi-beam environments. Meanwhile, DRL has proven effective in sequential decision-making tasks [11], yet its integration with lightweight positioning remains underexplored in NTNs.
Motivated by these gaps, this work extends the use of DRL beyond prior studies that focused primarily on handover optimization [5], load-balanced beam hopping [4], or cooperative scheduling [6]. We adopt the DQN framework [11], which approximates the action-value function with deep neural networks to enable efficient decision-making in large state-action spaces. Unlike previous DRL approaches requiring explicit CSI or heavy network coordination, the proposed framework applies DQN for beam positioning in multiple-input multiple-output (MIMO) LEO systems and integrates it with a WLS estimator [10]. This design allows adaptive beam selection and weighting directly from uplink pilot responses, eliminating CSI estimation, achieving robustness under interference-prone and GNSS-denied conditions, and significantly reducing computational complexity compared to conventional optimization or deep learning baselines. While CRLB-based optimization offers theoretical bounds, it is analytically intractable for fast-varying beam geometry. Hence, an augmented WLS estimator is adopted, whose efficiency asymptotically approaches the CRLB under Gaussian noise.
@@
  • A novel DRL-based beam positioning framework is proposed for MIMO LEO systems, extending the use of reinforcement learning beyond handover, beam-hopping, and cooperative scheduling to precise user positioning.
  • The proposed DQN dynamically selects and weights satellite beams, while a WLS estimator computes user positions, enabling robust localization without explicit CSI estimation.
  • The proposed framework substantially reduces computational complexity, enabling real-time deployment for resource-constrained LEO satellite platforms.
  • Simulation results show that the proposed method achieves more than 99.3% reduction in positioning error compared to baseline approaches, particularly under interference-prone and GNSS-denied scenarios.
@@ II System Model
II-A Coordinate System and Beam Geometry
We consider a LEO satellite constellation that continuously illuminates the Earth through multiple narrow beams. Within a short observation window, the ground-projected centers of visible beams are denoted by $\{\bm{c}_{i}\}_{i=1}^{M}$ in a local East-North-Up (ENU) or Earth-Centered Earth-Fixed (ECEF) coordinate system. The unknown user terminal (UT) position is represented by $\bm{x}\in\mathbb{R}^{d\times 1}$, where $d=2$ for planar and $d=3$ for full-space positioning.
For the $i$-th beam, the line-of-sight (LoS) unit vector from the beam center to the UT is defined as [12]
$\bm{h}_{i}\triangleq\frac{\bm{x}-\bm{c}_{i}}{\|\bm{x}-\bm{c}_{i}\|},\qquad i=1,\ldots,M,$ (1)
where $\bm{h}_{i}\in\mathbb{R}^{d\times 1}$ and $d=2$ or $3$ depending on the positioning dimension. The beam geometry $\{\bm{c}_{i},\bm{h}_{i}\}$ is assumed quasi-static within a 1–5 s window, updated by ephemeris and the scheduler.
@@ II-B Channel Model
Following the model in [6], the narrowband line-of-sight (LoS) channel between the satellite and the UT is dominated by a single deterministic path. Let $\bm{s}_{i}\in\mathbb{R}^{3}$ denote the satellite antenna phase-center position and $d_{i}=\|\bm{x}-\bm{s}_{i}\|$ the slant range between the satellite and the UT. Consider a uniform planar array (UPA) with $N_{x}\times N_{y}$ elements and total antenna count $N=N_{x}N_{y}$. The complex baseband channel vector of beam $i$ is expressed as
@@
where $\phi_{i}$ denotes the random phase uniformly distributed in $0,2\pi)$, $\bm{a}_{t}(\varphi_{i})$ is the uniform planar array (UPA) [13] transmit steering vector at the departure angle $\varphi_{i}$, and $\beta_{i}$ represents the large-scale channel gain defined as
@@
with $f_{c}$ denoting the carrier frequency, $c$ the speed of light, and $G_{t},G_{r}$ the transmit and receive antenna gains, respectively. The UPA steering vector is written as
$\bm{a}_{t}(\varphi_{i})=\bm{a}_{x}(\theta_{i}^{x})\otimes\bm{a}_{y}(\theta_{i}^{y}),$ (4)
where
$\bm{a}_{x}(\theta_{i}^{x})$ $=\frac{1}{\sqrt{N_{x}}}\!\left[1,e^{-j\pi\theta_{i}^{x}},\ldots,e^{-j\pi(N_{x}-1)\theta_{i}^{x}}\right]^{\!\top},$ (5)
$\bm{a}_{y}(\theta_{i}^{y})$ $=\frac{1}{\sqrt{N_{y}}}\!\left[1,e^{-j\pi\theta_{i}^{y}},\ldots,e^{-j\pi(N_{y}-1)\theta_{i}^{y}}\right]^{\!\top},$ (6)
and the direction cosines $(\theta_{i}^{x},\theta_{i}^{y})$ are determined by the azimuth and elevation of the UT relative to the satellite.
Given a transmit beamforming vector $\bm{f}_{i}\in\mathbb{C}^{N\times 1}$ for pilot transmission, the received scalar pilot signal and its corresponding SINR are given by
$y_{i}$ $=\bm{g}_{i}^{\!H}\bm{f}_{i}x_{i}+\sum_{k\neq i}\bm{g}_{i}^{\!H}\bm{f}_{k}x_{k}+v_{i},$ (7)
$\mathrm{SINR}_{i}$ $=\frac{|\bm{g}_{i}^{\!H}\bm{f}_{i}|^{2}}{\sum_{k\neq i}|\bm{g}_{i}^{\!H}\bm{f}_{k}|^{2}+\sigma^{2}},$ (8)
where $x_{i}$ is the unit-energy pilot symbol, $v_{i}\!\sim\!\mathcal{CN}(0,\sigma^{2})$ denotes complex Gaussian noise, and $\bm{f}_{k}\in\mathbb{C}^{N\times 1}$ represents the beamforming vector of the $k$-th co-channel beam $(k\!\neq\!i)$. The instantaneous per-beam link-quality metric $\mathrm{SNR}_{i}$ (or $\mathrm{SINR}_{i}$ under frequency reuse) computed from (2)-(8) is normalized to $[0,1]$ and later used as part of the DRL observation vector.
@@ II-C Observation Model
In conventional beam-intersection methods, the UT location is inferred from the overlap of beam coverage regions. To build a unified and analytically tractable model, we define a set of differentiable measurements $\bm{z}\in\mathbb{R}^{M}$ that depend on the geometric distances between $\bm{x}$ and the beam centers $\{\bm{c}_{i}\}$. For the $i$-th beam, a pseudorange-like measurement is modeled as
$z_{i}=\|\bm{x}-\bm{c}_{i}\|+b+n_{i},$ (9)
where $b$ represents the receiver clock bias and $n_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$ is zero-mean Gaussian noise whose variance reflects the effective SNR of beam $i$.
@@
$\bm{z}\approx\bm{H}\bm{x}+b\mathbbold{1}+\bm{n},$ (10)
where the $i$-th row of $\bm{H}\in\mathbb{R}^{M\times d}$ is $\bm{h}_{i}^{\top}$ defined in (1), and $\mathbbold{1}\in\mathbb{R}^{M\times 1}$.
@@ II-D Augmented Weighted Least Squares (WLS)
Let $\tilde{\bm{H}}=[\bm{H}\ \ \mathbbold{1}]\in\mathbb{R}^{M\times(d+1)}$ and $\tilde{\bm{x}}=[\bm{x}^{\top}\ b]^{\top}$. Given $\bm{W}=\mathrm{diag}(w_{1},\ldots,w_{M})$, the augmented WLS estimate is
@@ III Problem Formulation
From the system model above, the WLS estimator in (11) computes the UT position $\hat{\bm{x}}$ based on the beam-weight vector $\bm{w}=[w_{1},\ldots,w_{M}]^{\top}$. The instantaneous positioning error is defined as
@@
The objective of the beam positioning problem is to find the optimal weight vector $\bm{w}^{*}$ that minimizes the expected positioning error:
@@
The above optimization is nonlinear and highly dependent on the beam geometry through the matrix $\bm{H}$ in (10). Since $\bm{H}$ varies rapidly with satellite motion and beam scheduling, solving (13) analytically or by exhaustive search is computationally prohibitive. Conventional heuristic methods assign fixed or power-based weights that cannot adapt to dynamic link conditions.
To overcome these limitations, we reformulate the beam-weight optimization as a sequential decision-making problem, where an agent learns to adjust $\bm{w}$ based on observed beam features and positioning feedback. This motivates the proposed DRL approach detailed in Section IV. The continuous weight vector is discretized into a finite set of beam-selection and coarse-weight combinations to ensure tractable DQN training while maintaining sufficient action granularity.
IV DQN-WLS Based Beam Positioning
In this section, we present a DRL framework that integrates a DQN with a WLS estimator for accurate and adaptive beam-based positioning in LEO satellite systems. Unlike conventional geometry- or CSI-based methods, the proposed DQN-WLS scheme dynamically learns beam-weighting strategies directly from uplink pilot responses, thus eliminating explicit CSI estimation and reducing computational complexity.
@@ IV-A Markov Decision Process Formulation
The beam-weight optimization problem in (13) is reformulated as a Markov decision process (MDP) $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$. At each discrete time step $t$, the environment (BeamEnv) generates a state vector $\bm{s}_{t}\in\mathcal{S}$ that encapsulates both geometric and signal-level features of $M$ visible beams. These include normalized beam distances, azimuth-elevation indicators, per-beam SINR metrics, previous residual errors, and prior weight distributions.
Action parameterization. To comply with DQN’s discrete-action setting, we quantize the continuous beam weights into a finite action set constructed by top-$K$ beam selections and coarse weight levels $\{0,0.25,0.5,0.75,1\}$, followed by $\ell_{1}$ normalization. Near-duplicate actions are pruned by correlation screening to keep the action set tractable.
@@
$r_{t}=-\left(\frac{e_{t}}{\tau}\right)^{2}-\alpha\sum_{i=1}^{M}w_{t,i}(1-\mathrm{SINR}_{i})-\beta\big(1-\mathrm{Entropy}(\bm{w}_{t})\big),$ (14)
@@
which encourages exploration by promoting weight diversity. Here, $\tau$ controls the error-scaling factor, $\alpha$ penalizes low-quality beams, and $\beta$ regulates the entropy contribution. The discount factor $\gamma=0.99$ encourages long-term optimization over sequential interactions.
@@ IV-B Network Architecture
The DQN approximates the action-value function $Q_{\theta}(\bm{s},\bm{a})$ parameterized by $\theta$. The input layer concatenates all beam-level features with a total feature dimension $F=60$ for $M=10$ beams. Two fully connected hidden layers of 128 neurons each, activated by ReLU, are used to extract latent geometric representations. After pruning redundant combinations, the discrete action set contains approximately $10^{2}$ valid beam-weight configurations. The network outputs $|\mathcal{A}|$ Q-values over these quantized actions, and the selected action is mapped to a normalized weight vector $\bm{w}_{t}$.

### v3

@@ Abstract
This paper investigates a lightweight deep reinforcement learning (DRL)-assisted weighting framework for CSI-free multi-satellite positioning in LEO constellations, where each visible satellite provides one serving beam (one pilot response) per epoch. A discrete-action Deep Q-Network (DQN) learns satellite weights directly from received pilot measurements and geometric features, while an augmented weighted least squares (WLS) estimator provides physics-consistent localization and jointly estimates the receiver clock bias. The proposed hybrid design targets an accuracy-runtime trade-off rather than absolute supervised optimality. In a representative 2-D setting with 10 visible satellites, the proposed approach achieves sub-meter accuracy (0.395m RMSE) with low computational overhead, supporting practical deployment for resource-constrained LEO payloads.
@@ I Introduction
Despite these advances, most existing solutions still rely heavily on explicit channel state information (CSI) or involve computationally intensive optimization, limiting real-time applicability for resource-constrained LEO platforms. Traditional statistical estimation methods, such as weighted least squares (WLS) [10], widely used in positioning problems, provide reliable accuracy but lack adaptability in dynamic multi-satellite environments with heterogeneous link qualities. Meanwhile, DRL has proven effective in sequential decision-making tasks [11], yet its integration with lightweight positioning remains underexplored in NTNs.
Motivated by these gaps, this work extends the use of DRL beyond prior studies that focused primarily on handover optimization [5], load-balanced beam hopping [4], or cooperative scheduling [6]. We adopt the DQN framework [11], which approximates the action-value function with deep neural networks to enable efficient decision-making in large state-action spaces. Unlike previous DRL approaches requiring explicit CSI or heavy network coordination, the proposed framework applies DQN for multi-satellite positioning in multiple-input multiple-output (MIMO) LEO systems and integrates it with a WLS estimator [10]. This design allows adaptive beam selection and weighting directly from received pilot measurements. Thus, it eliminates explicit CSI estimation, achieving robustness under interference-prone and global navigation satellite system (GNSS)-denied conditions. In addition, it significantly reduces computational complexity compared to conventional optimization or deep learning baselines. While CRLB-based optimization offers theoretical bounds, it is analytically intractable for fast-varying satellite geometry. Hence, an augmented WLS estimator is adopted, whose efficiency asymptotically approaches the CRLB under Gaussian noise.
@@
  • CSI-free, lightweight hybrid design: We couple a discrete-action DQN satellite-weighting policy with an augmented WLS solver to enable CSI-free multi-satellite positioning, while maintaining a small on-board inference footprint.
  • Deployment-oriented action parameterization: We design a quantized satellite-weight action space compatible with practical on-board control granularity, improving numerical robustness of the downstream WLS inversion while remaining tractable for DQN training.
  • Accuracy-runtime evidence and insight: We provide a comparative evaluation highlighting the trade-off between positioning accuracy and computational cost across representative baselines, illustrating the regime where the proposed approach is attractive for real-time LEO operation.
@@ II System Model
Fig. 1 illustrates the proposed CSI-free multi-satellite positioning pipeline. At each epoch, $M$ visible LEO satellites are available, and each satellite schedules one serving beam to deliver a pilot observation. Pilot-derived features $\{y_{i}\}_{i=1}^{M}$, geometry information (e.g., $\{\bm{s}_{i},\bm{h}_{i}\}$), and the normalized residual Doppler magnitudes $\{|{\Delta f_{\mathrm{D}}}_{i}|T_{p}\}_{i=1}^{M}$ (with $T_{p}$ defined in (8)) constitute the DQN state. The DQN selects a discrete action that is mapped to a normalized satellite-weight vector, which is then applied to the augmented WLS estimator to jointly estimate the UT position and the receiver clock bias.
Figure 1: System overview of the proposed CSI-free DQN-WLS multi-satellite positioning framework.
II-A Coordinate System and Satellite geometry
We consider $M$ visible LEO satellites from a constellation. Let $\{\bm{s}_{i}\}_{i=1}^{M}$ denote the satellite positions in the earth-centered, earth-fixed (ECEF) coordinate system within a short observation window. The unknown UT position is denoted by $\bm{x}\in\mathbb{R}^{d\times 1}$.
For satellite $i$, the line-of-sight (LoS) unit vector from the satellite to the UT is
$\bm{h}_{i}\triangleq\frac{\bm{x}-\bm{s}_{i}}{\|\bm{x}-\bm{s}_{i}\|},\quad i=1,\ldots,M.$ (1)
where $\bm{h}_{i}\in\mathbb{R}^{d\times 1}$ and $d=2$ or $3$ depending on the positioning dimension. The satellite geometry $\{\bm{s}_{i},\bm{h}_{i}\}$ is assumed quasi-static within a 1 to 5 s window, updated by ephemeris and the scheduler.
@@ II-B Channel Model
Following the model in [6], the narrowband line-of-sight (LoS) channel between the satellite and the UT is dominated by a single deterministic path. Let $\bm{s}_{i}\in\mathbb{R}^{3}$ denote the satellite antenna phase-center position and $d_{i}=\|\bm{x}-\bm{s}_{i}\|$ the slant range between satellite $i$ and the UT. Consider a uniform planar array (UPA) [13] with $N_{x}\times N_{y}$ elements and total antenna count $N=N_{x}N_{y}$. The complex baseband channel vector from satellite $i$ is expressed as
@@
where $\phi_{i}$ denotes the random phase uniformly distributed in $0,2\pi)$, $\bm{a}_{t}(\varphi_{i})$ is the UPA transmit steering vector at the departure angle $\varphi_{i}$, and $\beta_{i}$ represents the large-scale channel gain defined as
@@
with $f_{c}$ denoting the carrier frequency, $c$ the speed of light, and $G_{t},G_{r}$ the transmit and receive antenna gains, respectively. The transmit steering vector $\bm{a}_{t}(\varphi_{i})$ follows a standard $N_{x}\times N_{y}$ UPA response model [13]. To reduce exposition, we omit its element-wise expression and directly use $\bm{a}_{t}(\varphi_{i})$ to compute the effective beamformed gain $|\bm{g}_{i}^{H}\bm{f}_{i}|^{2}$ in (5).
Within each epoch, the UT correlates $N_{p}$ received pilot samples to obtain one scalar pilot feature per satellite:
$y_{i}\triangleq\frac{1}{N_{p}}\sum_{n=0}^{N_{p}-1}r[n]\,x_{i}^{*},$ (4)
where $r[n]$ denotes the received pilot sample and $x_{i}$ is the known pilot symbol of satellite $i$. After ephemeris-based Doppler pre-compensation, the residual Doppler is modeled as ${\Delta f_{\mathrm{D}}}_{i}\sim\mathcal{N}(0,\sigma_{\Delta f_{\mathrm{D}}}^{2})$, i.i.d. across satellites and epochs.
Accordingly, we compute the instantaneous power-domain link-quality metric under frequency reuse as
$\mathrm{SINR}_{i}=\frac{|\bm{g}_{i}^{\!H}\bm{f}_{i}|^{2}}{\sum_{k\neq i}|\bm{g}_{k}^{\!H}\bm{f}_{k}|^{2}+\sigma^{2}},$ (5)
where the summation term accounts for frequency-reuse co-channel beams transmitted by other satellites on the same time-frequency resource, and $\bm{g}_{k}$ denotes the channel vector from satellite $k$ to the UT within the considered epoch. The residual Doppler impact is incorporated via an equivalent measurement-reliability penalty in (7)-(8). We emphasize that a full orthogonal frequency-division multiplexing (OFDM) inter-carrier interference (ICI) treatment is beyond the scope of this work and can be captured by an equivalent SINR/variance penalty when needed. The instantaneous per-satellite link-quality metric $\mathrm{SINR}_{i}$ under frequency reuse computed from (2)-(5) is normalized to $[0,1]$ and later used as part of the DRL observation vector.
@@ II-C Observation Model
In conventional multi-satellite positioning, the UT location is inferred from multiple range-like measurements collected from visible satellites. To build a unified and analytically tractable model, we define a set of differentiable measurements $\bm{z}\in\mathbb{R}^{M}$ that depend on the geometric distances between $\bm{x}$ and the satellite positions $\{\bm{s}_{i}\}$. For satellite $i$, a pseudorange-like measurement is modeled as
$z_{i}=\|\bm{x}-\bm{s}_{i}\|+b+n_{i},$ (6)
where $b$ is the receiver clock bias and $n_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})$ is zero-mean Gaussian noise whose variance reflects the effective link quality of satellite $i$. To incorporate residual Doppler, we model an equivalent reliability penalty on the range-like measurement. Specifically, we define an effective SINR as
$\mathrm{SINR}_{i}^{\mathrm{eff}}=\frac{\mathrm{SINR}_{i}}{1+\kappa\left(|{\Delta f_{\mathrm{D}}}_{i}|T_{p}\right)^{2}},$ (7)
where $\kappa$ is a scaling parameter controlling the Doppler-induced reliability penalty. The measurement noise variance is modeled as
$\sigma_{i}^{2}=\frac{\sigma_{0}^{2}}{\mathrm{SINR}_{i}^{\mathrm{eff}}},\qquad T_{p}=N_{p}T_{s},$ (8)
where $\sigma_{0}^{2}$ is a nominal noise level, $\kappa$ controls the Doppler-induced penalty strength, and $T_{p}$ denotes the pilot integration time used for pilot correlation/averaging within one epoch. This model captures that residual Doppler degrades pilot correlation quality and thus reduces the reliability of range-like measurements, which is directly reflected in the WLS weighting.
@@
$\bm{z}\approx\bm{H}\bm{x}+b\mathds{1}+\bm{n},$ (9)
where the $i$-th row of $\bm{H}\in\mathbb{R}^{M\times d}$ is $\bm{h}_{i}^{\top}$ defined in (1), $\mathds{1}\in\mathbb{R}^{M\times 1}$, and $\bm{x}_{0}$ is an initial position estimate.
@@ II-D Augmented Weighted Least Squares (WLS)
Let $\tilde{\bm{H}}=[\bm{H}\ \ \mathds{1}]\in\mathbb{R}^{M\times(d+1)}$ and $\tilde{\bm{x}}=[\bm{x}^{\top}\ b]^{\top}$. Given $\bm{W}=\mathrm{diag}(w_{1},\ldots,w_{M})$, the augmented WLS estimate is
@@ III Problem Formulation
From the system model above, the WLS estimator in (10) computes the UT position $\hat{\bm{x}}$ based on the satellite weight vector $\bm{w}=[w_{1},\ldots,w_{M}]^{\top}$. The instantaneous positioning error is defined as
@@
The objective of the multi-satellite positioning problem is to find the optimal weight vector $\bm{w}^{*}$ that minimizes the expected positioning error:
@@
The above optimization is nonlinear and highly dependent on the satellite geometry through the matrix $\bm{H}$ in (9). Since $\bm{H}$ varies rapidly with satellite motion and the time-varying set of visible/serving satellite links, solving (12) analytically or by exhaustive search is computationally prohibitive. Conventional heuristic methods assign fixed or power-based weights that cannot adapt to dynamic link conditions.
To overcome these limitations, we reformulate the satellite-weight optimization as a sequential decision-making problem, where an agent learns to adjust $\bm{w}$ based on observed satellite-link features and positioning feedback. This motivates the proposed DRL approach detailed in Section IV. To support practical deployability and numerical stability, the continuous weight vector is quantized into a finite action set: (i) it matches hardware-friendly weight granularity (e.g., codebook-like control), (ii) it mitigates ill-conditioning in the augmented WLS inversion under noisy beams, and (iii) it yields a tractable discrete action space for DQN training.
IV DQN-WLS Based Multi-Satellite Positioning
In this section, we present a DRL framework that integrates a DQN with a WLS estimator for accurate and adaptive multi-satellite positioning in LEO satellite systems. Unlike conventional geometry- or CSI-based methods, the proposed DQN-WLS scheme learns satellite-weighting strategies directly from received pilot measurements and geometric features, thus eliminating explicit CSI estimation and reducing computational complexity.
@@ IV-A Markov Decision Process Formulation
The beam-weight optimization problem in (12) is reformulated as a Markov decision process (MDP) $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$. At each discrete time step $t$, the environment (BeamEnv) generates a state vector $\bm{s}_{t}\in\mathcal{S}$ that encapsulates both geometric and signal-level features of $M$ visible satellite links. These include normalized satellite distances (or slant ranges), azimuth-elevation indicators, per-satellite SINR metrics, previous residual errors, and prior weight distributions.
Action parameterization. To comply with DQN’s discrete-action setting, we quantize the continuous satellite weights into a finite action set constructed by top-$K$ satellite selections and coarse weight levels $\{0,0.25,0.5,0.75,1\}$, followed by $\ell_{1}$ normalization. Near-duplicate actions are pruned by correlation screening to keep the action set tractable.
@@
$r_{t}=-\left(\frac{e_{t}}{\tau}\right)^{2}-\alpha\sum_{i=1}^{M}w_{t,i}\big(1-\mathrm{SINR}_{i}^{\mathrm{eff}}\big)-\beta\big(1-\mathrm{Entropy}(\bm{w}_{t})\big),$ (13)
@@
which encourages exploration by promoting weight diversity. Here, $\tau$ controls the error-scaling factor, $\alpha$ penalizes low-quality satellite links, and $\beta$ regulates the entropy contribution. The discount factor $\gamma=0.99$ encourages long-term optimization over sequential interactions.
@@ IV-B Network Architecture
The DQN approximates the action-value function $Q_{\theta}(\bm{s},\bm{a})$ parameterized by $\theta$. The input layer concatenates per-satellite features with a total feature dimension $F=60$ for $M=10$ satellites. Two fully connected hidden layers of 128 neurons each, activated by ReLU, are used to extract latent geometric representations. After pruning redundant combinations, the discrete action set contains approximately $10^{2}$ valid satellite-weight configurations. The network outputs $|\mathcal{A}|$ Q-values over these quantized actions, and the selected action is mapped to a normalized weight vector $\bm{w}_{t}$.